rock was shot horizontally with a speed of 15 m/s, from the roof of a building. It

landed on the ground 100 m away from that building.

a. Calculate the time taken by the rock to reach its final position.


1
Expert's answer
2021-02-19T10:28:16-0500

a)

"x=v_0t,""t=\dfrac{x}{v_0}=\dfrac{100\ m}{15\ \dfrac{m}{s}}=6.67\ s."